# Fast Estimation of Wasserstein Distances via Regression on Sliced Wasserstein Distances

### v1

@@ Abstract
We address the problem of efficiently computing Wasserstein distances for multiple pairs of distributions drawn from a meta-distribution. To this end, we propose a fast estimation method based on regressing Wasserstein distance on sliced Wasserstein (SW) distances. Specifically, we leverage both standard SW distances, which provide lower bounds, and lifted SW distances, which provide upper bounds, as predictors of the true Wasserstein distance. To ensure parsimony, we introduce two linear models: an unconstrained model with a closed-form least-squares solution, and a constrained model that uses only half as many parameters. We show that accurate models can be learned from a small number of distribution pairs. Once estimated, the model can predict the Wasserstein distance for any pair of distributions via a linear combination of SW distances, making it highly efficient. Empirically, we validate our approach on diverse tasks, including Gaussian mixtures, point-cloud classification, and Wasserstein-space visualizations for 3D point clouds. Across various datasets such as MNIST point clouds, ShapeNetV2, MERFISH Cell Niches, and scRNA-seq, our method consistently provides a better approximation of Wasserstein distance than the state-of-the-art Wasserstein embedding model, Wasserstein Wormhole, particularly in low-data regimes. Finally, we demonstrate that our estimator can also accelerate Wormhole training, yielding RG-Wormhole.
footnotetext: * Equal Contribution
@@ 1 Introduction
Optimal Transport (OT) and Wasserstein distances villani2009optimal ; peyre2019computational have become essential tools in machine learning, widely used for quantifying the similarity or dissimilarity between probability distributions. Fundamentally, the Wasserstein distance measures the minimum cost required to "transport" mass from one distribution to another, effectively capturing the underlying geometry of the data. Thanks to their clear geometric interpretation and mathematical robustness, Wasserstein distances have found applications across various fields, such as generative modeling genevay2018learning , computational biology bunne2023learning , chemistry wu2023improving , and image processing feydy2017optimal . Despite its utility, computing the exact Wasserstein distance is computationally expensive. It typically requires solving a large-scale linear program to find an optimal transport plan, with a time complexity of $\mathcal{O}(n^{3}\log n)$ for discrete distributions of size $n$. This high cost severely limits its use in large-scale or real-time settings.
@@
In many applications, Wasserstein distances are computed (repeatedly) for many pairs of distributions, e.g., dataset comparisons alvarez2020geometric , 3D point-cloud autoencoder achlioptas2018learning , point-cloud nearest neighbor classification/regression rubner1998metric , learning embeddings for distributions kolouri2021wasserstein , density-density regression chen2023wasserstein , and so on. Therefore, the high computational complexities of the Wasserstein distance become the main bottleneck to scaling up these applications. As a result, speeding up the computation of the Wasserstein distance has become a vital task in practice.
@@
To address this bottleneck, a straightforward improvement is to speed up the computation of the Wasserstein distance. For example, entropic regularization cuturi2013sinkhorn enables fast approximation via Sinkhorn iterations, while other methods exploit the structure in the transport plan, such as low-rank approximations scetbon2021low . In addition, some approaches rely on strong structural assumptions, such as the Bures-Wasserstein metric dowson1982frechet gives a closed-form solution for the exact 2-Wasserstein distance ($W_{2}$) under the Gaussian assumption on distributions.
@@
Another approach is to cast computing Wasserstein distances for many pairs of distributions as a learning problem, i.e., learning a model first to predict the Wasserstein distance given any pair of distributions, then use the model later for the mentioned downstream tasks. For example, Deep Wasserstein Embedding (DWE) courty2018learning trains a Siamese convolutional network to match OT distances between 2D images, while Wasserstein Wormhole haviv2024wasserstein employs transformer-based architectures to learn embeddings of distributions, allowing Euclidean distances in the learned space to approximate Wasserstein distances efficiently. While effective, these deep learning-based methods require significant computational resources and time to train, and their performance may degrade when limited training data are available. Moreover, these approaches are limited to empirical distributions because of the use of neural networks.
@@
In this work, we propose a novel approach to predict the Wasserstein distance without relying on any neural networks or learned embeddings. Moreover, the proposed approach relies on a parsimonious model and can handle both continuous and discrete distributions. In particular, we propose to regress the Wasserstein distance on sliced Wasserstein (SW) distances rabin2010regularization ; maheyfast2023 ; nguyen2023energy ; liu2025expected ; deshpande2019max ; rowland2019orthogonal . In greater detail, we introduce linear models with Wasserstein distances as the response and SW distances as the predictors. We provide estimates of the models via efficient least-squares estimates. In addition, since sliced Wasserstein distances have low computational complexity, the resulting Wasserstein regressor is computationally efficient.
@@
1. We introduce the first regression framework where the Wasserstein distance serves as the response variable and various sliced Wasserstein (SW) distances act as predictors, in the setting of random pairs of distributions. This framework not only uncovers the relationship between the Wasserstein distance and its SW-based approximations but also enables efficient estimation of the Wasserstein distance. Specifically, we use SW distance bonneel2015sliced , Max-SW deshpande2019max , and energy-based SW nguyen2023energy , all of which provide lower bounds on the Wasserstein distance, as predictors. In addition, we incorporate lifted SW distances, which provide upper bounds, including projected Wasserstein rowland2019orthogonal , Minimum SW generalized geodesics maheyfast2023 , and expected sliced distance liu2025expected .
@@ 2 Preliminaries
Wasserstein distance. Wasserstein-$p$ ($p\geq 1$) distance villani2008optimal ; peyre2020computational between two distributions $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ (dimension $d\geq 1$) is defined as:
@@
where $\Pi(\mu,\nu)=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})\}\mid\int_{\mathbb{R}^{d}}d\pi(x,y)=\mu(x),\int_{\mathbb{R}^{d}}d\pi(x,y)=\nu(y)\right\}$ is the set of all transportation plans i.e., joint distributions which have marginals be two comparing distributions. When $\mu$ and $\nu$ are discrete distributions i.e., $\mu=\sum_{i=1}^{n}\alpha_{i}\delta_{x_{i}}$ ($n\geq 1$) and $\nu=\sum_{j=1}^{m}\beta_{j}\delta_{y_{j}}$ ($m\geq 1$) where $\sum_{i=1}^{N}\alpha_{i}=\sum_{j=1}^{m}\beta_{j}=1$ and $\alpha_{i}\geq 0,\beta_{j}\geq 0$ for all $i=1,\ldots,n$ and $j=1,\ldots,m$, Wasserstein distance between $\mu$ and $\nu$ defined as: $W_{p}^{p}(\mu,\nu)=\min_{\gamma\in\Gamma(\alpha,\beta)}\sum_{i=1}^{n}\sum_{j=1}^{m}\|x_{i}-y_{j}\|_{p}^{p}\gamma_{ij},$ where $\Gamma(\alpha,\beta)=\{\gamma\in\mathbb{R}^{n\times m}_{+}\mid\gamma\mathbf{1}=\alpha,\gamma^{\top}\mathbf{1}=\beta\}$. Without loss of generality, we assume that $n\geq m$. Therefore, the time complexity for solving this linear programming is $\mathcal{O}(n^{3}\log n)$ peyre2019computational and $\mathcal{O}(n^{2})$, which are expensive.
@@
$F_{\mu}^{-1}(t)=\sum_{i=1}^{n}x_{(i)}I\left(\sum_{j=1}^{i-1}\alpha_{(j)}\leq t\leq\sum_{j=1}^{i}\alpha_{(j)}\right),F_{\nu}^{-1}(t)=\sum_{j=1}^{m}y_{(j)}I\left(\sum_{i=1}^{j-1}\beta_{(i)}\leq t\leq\sum_{i=1}^{j}\beta_{(i)}\right),$
@@
Random Projection. A key technique that plays a vital role in later discussion is random projection. We consider a function $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ where $\theta\sim\sigma(\theta)$ ($\sigma(\theta)\sim\mathcal{P}(\mathbb{S}^{d-1})$) is a random variable. For simplicity, we consider the traditional setup where $\theta\sim\mathcal{U}(\mathbb{S}^{d-1})$ and $P_{\theta}(x)=\langle\theta,x\rangle$ bonneel2015sliced ; rabin2012wasserstein . However, the following discussion holds for any other types of projections kolouri2019generalized ; bonet2023hyperbolic ; bonet2024sliced ; bonet2023sliced . For $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, one-dimensional projected Wasserstein distance with $P_{\theta}$ is defined as:
@@
The second approach to construct a Wasserstein-type discrepancy from one-dimensional projection is using lifted transportation plan. There are many ways to construct such lifted plan using disintegration of measures muzellec2019subspace ; tanguy2025sliced . In practice, the most used way liu2025expected ; tanguy2025sliced is:
@@
where $\pi^{\theta}\in\Pi(\mu,\nu)$ is the lifted transportation plan, $\pi_{(P_{\theta}\sharp\mu,P_{\theta}\sharp\nu)}$ is the optimal transport plan between $P_{\theta}\sharp\mu$ and $P_{\theta}\sharp\nu$, $\mu_{t_{1}}$ and $\nu_{t_{2}}$ are disintegration of $\mu$ and $\nu$ at $t_{1}$ and $t_{2}$ the function $P_{\theta}$, and $\otimes$ denotes the product of measures. When dealing with discrete measures $\mu$ and $\nu$, $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ can still be computed efficiently maheyfast2023 ; liu2025expected i.e., $\mathcal{O}(n\log n)$ in time and $\mathcal{O}(n)$ in space (assumed that $n>m$). The quantity $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is known as lifted cost tanguy2025sliced or sliced Wasserstein generalized geodesic maheyfast2023 ; liu2025expected . From previous work nguyen2023energy ; maheyfast2023 ; tanguy2023convergence , we know the following relationship $\underline{W}_{p}(\mu,\nu;P_{\theta})\leq W_{p}(\mu,\nu)\leq\overline{W}_{p}(\mu,\nu;P_{\theta})$.
@@ 3.1 Sliced Wasserstein and Lifted Sliced Wasserstein
Sliced Wasserstein distances. Given $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, a sliced Wasserstein-$p$ distance can be defined as follows rabin2012wasserstein ; nguyen2025introduction :
@@
where $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ is the projection function, $\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is the one-dimensional projected Wasserstein distance (equation 2), and $\sigma\in\mathcal{P}(\mathbb{S}^{d-1})$ is the slicing distribution. By changing the slicing distribution, we can obtain variants of SW. There are three main ways: 1. Fixed prior: The simplest way is to choose $\sigma$ to be a fixed and known distribution, e.g., a uniform distribution $\mathcal{U}(\mathbb{S}^{d-1})$ as in the conventional SW rabin2012wasserstein . 2. Optimization-based: We can also find $\sigma$ that prioritizes some realizations of $\theta$ that satisfies a notion of informativeness. For example, $\sigma$ can put more masses to realizations of $\theta$ where $\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ have high value, i.e., setting informativeness as discriminativeness. For example, we can find $\sigma$ by solving nguyen2021distributional : $\sup_{\sigma\in\mathcal{M}(\mathbb{S}^{d-1})}\mathbb{E}_{\theta\sim\sigma}[\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})],$ where $\mathcal{M}(\mathbb{S}^{d-1})\subset\mathcal{P}(\mathbb{S}^{d-1})$ be a set of probability measures on $\mathbb{S}^{d-1}$. When $\mathcal{M}(\mathbb{S}^{d-1})=\{\delta_{\theta}\mid\theta\in\mathbb{S}^{d-1}\}$, max sliced Wasserstein distance deshpande2019max is obtained: $\text{Max-}SW(\mu,\nu)=\max_{\theta\in\mathbb{S}^{d-1}}\underline{W}_{p}(\mu,\nu;P_{\theta})].$ 3. Energy-based: An optimization-free way to select $\sigma$ is to design it as an energy-based distribution with the unnormalized density: $p_{\sigma}(\theta)\propto f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})),$ where $f$ is often chosen to be an increasing function on the positive real line, i.e., an exponential function. This choice of slicing distribution leads to energy-based SW (EBSW) nguyen2023energy .
@@
Empirical estimation. For SW, Monte Carlo estimation is used to approximate the distance: $\widehat{SW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\frac{1}{L}\sum_{l=1}^{L}\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$ where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$ ($L>0$) are projecting directions (other sampling techniques can also be used nguyen2024quasimonte ; nguyen2024control ; sisouk2025a ). For Max-SW, we can use $\hat{\theta}_{T}$ which is the solution of an optimization algorithm with $T>0$ iterations, e.g., projected gradient ascent nietert2022statistical or Riemannian gradient ascent lin2020projection : $\widehat{\text{Max-}SW}_{p}^{p}(\mu,\nu;\hat{\theta}_{T})=\underline{W}_{p}^{p}(\mu,\nu;P_{\hat{\theta}_{T}}).$ For EBSW, one simple way to estimate the distance is to use importance sampling: $\widehat{EBSW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\sum_{l=1}^{L}\hat{w}_{l}\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$ where $\hat{w}_{l}=\frac{f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}))}{\sum_{l^{\prime}=1}^{L}f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l^{\prime}}})}$ and $\theta_{1},\ldots,\theta_{L}\sim\mathcal{U}(\mathbb{S}^{d-1})$.
@@
Lower bounds. We summarize the connection between SW, Max-SW, EBSW, and Wasserstein distance in the following remark. The detail of the proof can be found in nguyen2023energy .
@@
Lifted sliced Wasserstein distances. Given $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, a lifted sliced Wasserstein-$p$ distance can be defined as follows rowland2019orthogonal :

### v2

@@ Abstract
We address the problem of efficiently computing Wasserstein distances for multiple pairs of distributions drawn from a meta-distribution. To this end, we propose a fast estimation method based on regressing Wasserstein distance on sliced Wasserstein (SW) distances. Specifically, we leverage both standard SW distances, which provide lower bounds, and lifted SW distances, which provide upper bounds, as predictors of the true Wasserstein distance. To ensure parsimony, we introduce two linear models: an unconstrained model with a closed-form least-squares solution, and a constrained model that uses only half as many parameters. We show that accurate models can be learned from a small number of distribution pairs. Once estimated, the model can predict the Wasserstein distance for any pair of distributions via a linear combination of SW distances, making it highly efficient. Empirically, we validate our approach on diverse tasks, including Gaussian mixtures, point-cloud classification, and Wasserstein-space visualizations for 3D point clouds. Across various datasets such as MNIST point clouds, ShapeNetV2, MERFISH Cell Niches, and scRNA-seq, our method consistently provides a better approximation of Wasserstein than the state-of-the-art method, Wasserstein Wormhole, and classical methods, particularly in low-data regimes. To illustrate its robustness, we also experiment the method with intra- and inter-class settings. Finally, we demonstrate that RG can accelerate Wasserstein Wormhole training, yielding RG-Wormhole.  [^1] Code is published at https://github.com/hainn2803/Regression-Wasserstein.
@@ 1 Introduction
Optimal Transport (OT) and Wasserstein distances (Villani, 2009; Peyré & Cuturi, 2019) have become essential tools in machine learning, widely used for quantifying the similarity or dissimilarity between probability distributions. Fundamentally, the Wasserstein distance measures the minimum cost required to "transport" mass from one distribution to another, effectively capturing the underlying geometry of the data. Thanks to their clear geometric interpretation and mathematical robustness, Wasserstein distances have found applications across various fields, such as generative modeling Genevay et al. (2018), computational biology Bunne et al. (2023), chemistry Wu et al. (2023), and image processing Feydy et al. (2017). Despite its utility, computing the exact Wasserstein distance is computationally expensive. It typically requires solving a large-scale linear program to find an optimal transport plan, with a time complexity of $\mathcal{O}(n^{3}\log n)$ for discrete distributions of size $n$. This high cost severely limits its use in large-scale or real-time settings.
@@
In many applications, Wasserstein distances are computed (repeatedly) for many pairs of distributions instead of a single pair, e.g., dataset comparisons (Alvarez-Melis & Fusi, 2020), 3D point-cloud autoencoder (Achlioptas et al., 2018), point-cloud nearest neighbor classification/regression (Rubner et al., 1998), learning embeddings for distributions (Kolouri et al., 2021), density-density regression (Chen et al., 2023), and so on. Therefore, the high computational complexities of the Wasserstein distance become the main bottleneck to scaling up these applications. As a result, speeding up the computation of the Wasserstein distance has become a vital task in practice.
@@
To address this bottleneck, a straightforward improvement is to speed up the computation of the Wasserstein distance. For example, entropic regularization (Cuturi, 2013) enables fast approximation via Sinkhorn iterations, while other methods exploit the structure in the transport plan, such as low-rank approximations (Scetbon et al., 2021). In addition, some approaches rely on strong structural assumptions, such as the Bures-Wasserstein metric (Dowson & Landau, 1982) gives a closed-form solution for the exact 2-Wasserstein distance ($W_{2}$) under the Gaussian assumption on distributions.
@@
Another approach is to cast computing Wasserstein distances for many pairs of distributions as a learning problem, i.e., learning a model first to predict the Wasserstein distance given any pair of distributions, then use the model later for the mentioned downstream tasks. For example, Deep Wasserstein Embedding (DWE) (Courty et al., 2018) trains a Siamese convolutional network to match OT distances between 2D images, while Wasserstein Wormhole (Haviv et al., 2024) employs transformer-based architectures to learn embeddings of distributions, allowing Euclidean distances in the learned space to approximate Wasserstein distances efficiently. While effective, these deep learning-based methods require significant computational resources and time to train, and their performance may degrade when limited training data are available. Moreover, these approaches are limited to empirical distributions because of the use of neural networks.
@@
In this work, we propose a novel approach to predict the Wasserstein distance without relying on any neural networks or learned embeddings. Moreover, the proposed approach relies on a parsimonious model and can handle both continuous and discrete distributions. In particular, we propose to regress the Wasserstein distance on sliced Wasserstein (SW) distances (Rabin et al., 2010; Mahey et al., 2023; Nguyen & Ho, 2023; Liu et al., 2025; Deshpande et al., 2019; Rowland et al., 2019). In greater detail, we introduce linear models with Wasserstein distances as the response and SW distances as the predictors. We provide estimates of the models via efficient least-squares estimates. In addition, since sliced Wasserstein distances have low computational complexity, the resulting Wasserstein regressor is computationally efficient.
@@
1. We introduce the first regression framework where the Wasserstein distance serves as the response variable and various sliced Wasserstein (SW) distances act as predictors, in the setting of random pairs of distributions. This framework not only uncovers the relationship between the Wasserstein distance and its SW-based approximations but also enables efficient estimation of the Wasserstein distance. Specifically, we use SW distance (Bonneel et al., 2015), Max-SW (Deshpande et al., 2019), and energy-based SW (Nguyen & Ho, 2023), all of which provide lower bounds on the Wasserstein distance, as predictors. In addition, we incorporate lifted SW distances, which provide upper bounds, including projected Wasserstein (Rowland et al., 2019), minimum SW generalized geodesics (Mahey et al., 2023), and expected sliced distance (Liu et al., 2025).
@@ 2 Preliminaries
Wasserstein distance. Wasserstein-$p$ ($p\geq 1$) distance Villani (2008); Peyré et al. (2019) between two distributions $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ (dimension $d\geq 1$) is defined as:
@@
where $\Pi(\mu,\nu)=\left\{\pi\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})\}\mid\int_{\mathbb{R}^{d}}d\pi(x,y)=\mu(x),\int_{\mathbb{R}^{d}}d\pi(x,y)=\nu(y)\right\}$ is the set of all transportation plans i.e., joint distributions which have marginals be two comparing distributions. When $\mu$ and $\nu$ are discrete distributions i.e., $\mu=\sum_{i=1}^{n}\alpha_{i}\delta_{x_{i}}$ ($n\geq 1$) and $\nu=\sum_{j=1}^{m}\beta_{j}\delta_{y_{j}}$ ($m\geq 1$) where $\sum_{i=1}^{N}\alpha_{i}=\sum_{j=1}^{m}\beta_{j}=1$ and $\alpha_{i}\geq 0,\beta_{j}\geq 0$ for all $i=1,\ldots,n$ and $j=1,\ldots,m$, Wasserstein distance between $\mu$ and $\nu$ defined as:
$W_{p}^{p}(\mu,\nu)=\min_{\gamma\in\Gamma(\alpha,\beta)}\sum_{i=1}^{n}\sum_{j=1}^{m}\|x_{i}-y_{j}\|_{p}^{p}\gamma_{ij},$
where $\Gamma(\alpha,\beta)=\{\gamma\in\mathbb{R}^{n\times m}_{+}\mid\gamma\mathbf{1}=\alpha,\gamma^{\top}\mathbf{1}=\beta\}$. Without loss of generality, we assume that $n\geq m$. Therefore, the time complexity for solving this linear programming is $\mathcal{O}(n^{3}\log n)$ Peyré & Cuturi (2019) and $\mathcal{O}(n^{2})$, which are expensive.
@@
$F_{\mu}^{-1}(t)=\sum_{i=1}^{n}x_{(i)}I\left(\sum_{j=1}^{i-1}\alpha_{(j)}<t\leq\sum_{j=1}^{i}\alpha_{(j)}\right),F_{\nu}^{-1}(t)=\sum_{j=1}^{m}y_{(j)}I\left(\sum_{i=1}^{j-1}\beta_{(i)}<t\leq\sum_{i=1}^{j}\beta_{(i)}\right),$
@@
Random Projection. A key technique that plays a vital role in later discussion is random projection. We consider a function $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ where $\theta\sim\sigma(\theta)$ ($\sigma(\theta)\sim\mathcal{P}(\mathbb{S}^{d-1})$) is a random variable. For simplicity, we consider the traditional setup where $\theta\sim\mathcal{U}(\mathbb{S}^{d-1})$ and $P_{\theta}(x)=\langle\theta,x\rangle$ (Bonneel et al., 2015; Rabin et al., 2012). However, the following discussion holds for any other types of projections (Kolouri et al., 2019; Bonet et al., 2023b; 2025; c). For $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, one-dimensional projected Wasserstein distance with $P_{\theta}$ is defined as:
@@
The second approach to construct a Wasserstein-type discrepancy from one-dimensional projection is using lifted transportation plan. There are many ways to construct such lifted plan using disintegration of measures (Muzellec & Cuturi, 2019; Tanguy et al., 2025). In practice, the most used way (Liu et al., 2025; Tanguy et al., 2025) is:
@@
where $\pi^{\theta}\in\Pi(\mu,\nu)$ is the lifted transportation plan, $\pi_{(P_{\theta}\sharp\mu,P_{\theta}\sharp\nu)}$ is the optimal transport plan between $P_{\theta}\sharp\mu$ and $P_{\theta}\sharp\nu$, $\mu_{t_{1}}$ and $\nu_{t_{2}}$ are disintegration of $\mu$ and $\nu$ at $t_{1}$ and $t_{2}$ the function $P_{\theta}$, and $\otimes$ denotes the product of measures. When dealing with discrete measures $\mu$ and $\nu$, $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ can still be computed efficiently (Mahey et al., 2023; Liu et al., 2025) i.e., $\mathcal{O}(n\log n)$ in time and $\mathcal{O}(n)$ in space (assumed that $n>m$). The quantity $\overline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is known as lifted cost (Tanguy et al., 2025) or sliced Wasserstein generalized geodesic (Mahey et al., 2023; Liu et al., 2025). From previous work (Nguyen & Ho, 2023; Mahey et al., 2023; Tanguy, 2023), we know the following relationship $\underline{W}_{p}(\mu,\nu;P_{\theta})\leq W_{p}(\mu,\nu)\leq\overline{W}_{p}(\mu,\nu;P_{\theta})$.
@@ 3.1 Sliced Wasserstein and Lifted Sliced Wasserstein
Sliced Wasserstein distances. Given $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, a sliced Wasserstein-$p$ distance can be defined as follows (Rabin et al., 2012; Nguyen, 2025):
@@
where $P_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ is the projection function, $\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ is the one-dimensional projected Wasserstein distance (equation 2), and $\sigma\in\mathcal{P}(\mathbb{S}^{d-1})$ is the slicing distribution. By changing the slicing distribution, we can obtain variants of SW. There are three main ways:
1. Fixed prior: The simplest way is to choose $\sigma$ to be a fixed and known distribution, e.g., the uniform distribution $\mathcal{U}(\mathbb{S}^{d-1})$ as in the conventional SW (Rabin et al., 2012).
2. Optimization-based: We can also find $\sigma$ that prioritizes some realizations of $\theta$ that satisfies a notion of informativeness. For example, $\sigma$ can put more masses to realizations of $\theta$ where $\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})$ have high value, i.e., setting informativeness as discriminativeness. For example, we can find $\sigma$ by solving (Nguyen et al., 2021):
$\sup_{\sigma\in\mathcal{M}(\mathbb{S}^{d-1})}\mathbb{E}_{\theta\sim\sigma}[\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})],$
where $\mathcal{M}(\mathbb{S}^{d-1})\subset\mathcal{P}(\mathbb{S}^{d-1})$ be a set of probability measures on $\mathbb{S}^{d-1}$. When $\mathcal{M}(\mathbb{S}^{d-1})=\{\delta_{\theta}\mid\theta\in\mathbb{S}^{d-1}\}$, max sliced Wasserstein distance (Deshpande et al., 2019) is obtained: $\text{Max-}SW(\mu,\nu)=\max_{\theta\in\mathbb{S}^{d-1}}\underline{W}_{p}(\mu,\nu;P_{\theta})].$
3. Energy-based: An optimization-free way to select $\sigma$ is to design it as an energy-based distribution with the unnormalized density: $p_{\sigma}(\theta)\propto f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta})),$ where $f$ is often chosen to be an increasing function on the positive real line, i.e., an exponential function. This choice of slicing distribution leads to energy-based SW (EBSW) (Nguyen & Ho, 2023).
Empirical estimation. For SW, Monte Carlo estimation is used to approximate the distance:
$\widehat{SW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\frac{1}{L}\sum_{l=1}^{L}\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$
@@
where $\theta_{1},\ldots,\theta_{L}\overset{i.i.d}{\sim}\mathcal{U}(\mathbb{S}^{d-1})$ ($L>0$) are projecting directions (other sampling techniques can also be used (Nguyen et al., 2024; Nguyen & Ho, 2024; Sisouk et al., 2025)). For Max-SW, we can use $\hat{\theta}_{T}$ which is the solution of an optimization algorithm with $T>0$ iterations, e.g., projected gradient ascent (Nietert et al., 2022) or Riemannian gradient ascent Lin et al. (2020): $\widehat{\text{Max-}SW}_{p}^{p}(\mu,\nu;\hat{\theta}_{T})=\underline{W}_{p}^{p}(\mu,\nu;P_{\hat{\theta}_{T}}).$ For EBSW, one simple way to estimate the distance is to use importance sampling:
$\widehat{EBSW}_{p}^{p}(\mu,\nu;\theta_{1},\ldots,\theta_{L})=\sum_{l=1}^{L}\hat{w}_{l}\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}),$
where $\hat{w}_{l}=\frac{f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l}}))}{\sum_{l^{\prime}=1}^{L}f(\underline{W}_{p}^{p}(\mu,\nu;P_{\theta_{l^{\prime}}})}$ and $\theta_{1},\ldots,\theta_{L}\sim\mathcal{U}(\mathbb{S}^{d-1})$.
@@
Lower bounds. We summarize the connection between SW, Max-SW, EBSW, and Wasserstein distance in the following remark. The detail of the proof can be found in Nguyen & Ho (2023).
@@
Lifted sliced Wasserstein distances. Given $\mu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, a lifted sliced Wasserstein-$p$ distance can be defined as follows (Rowland et al., 2019):